# ShiftAddViT: Mixture of Multiplication Primitives Towards Efficient Vision Transformers

### v4

@@ 1 Introduction
Vision Transformers (ViTs) have emerged as powerful vision backbone substitutes to convolutional neural networks (CNNs) due to their impressive performance transformer ; vit . However, ViTs’ state-of-the-art (SOTA) accuracy comes at the price of prohibitive hardware latency and energy consumption for both training and inference, limiting their prevailing deployment and applications liu2022ecoformer ; shu2021adder , even though pre-trained ViT models are publicly available. The bottlenecks are twofold based on many profiling results: attention and MLPs, e.g., they account for 36%/63% of the FLOPs and 45%/46% of the latency when running DeiT-Base DeiT on an NVIDIA Jetson TX2 GPU you2022vitcod ; fan2022adaptable ). Previous efficient ViT solutions mostly focus on macro-architecture designs graham2021levit ; mehta2021mobilevit ; li2022mvitv2 ; DeiT ; chen2021autoformer ; wang2022pvt and linear attention optimization wang2020linformer ; shen2021efficient ; cai2022efficientvit ; choromanski2021rethinking ; chen2021scatterbrain ; you2023castling . However, they pay less attention on reducing the underlying dominant multiplications as well as on co-optimizing both attention and MLPs.
We identify that one of the most effective yet still missing opportunities for accelerating ViTs is to reparameterize their redundant multiplications with a mixture of multiplication primitives, i.e., bitwise shifts and adds. The idea is drawn from the common hardware design practice in computer architecture or digital signal processing. That is, the multiplication can be replaced by bitwise shifts and adds xue1986adaptive ; gwee2008low . Such a hardware-inspired approach can lead to an efficient and fast hardware implementation without compromising the model accuracy. As such, this paper embodies the shift&add idea in ViTs towards a new type of multiplication-reduced model. Moreover, unlike previous ShiftAddNet you2020shiftaddnet which requires full and slow training and dedicated hardware accelerator support, this work starts from pre-trained ViTs to avoid the tremendous cost of training from scratch and targets end-to-end inference speedups, i.e., accelerating both attention and MLPs, on GPUs.
@@
The aforementioned shift&add concept uniquely inspires our design of naturally hardware-friendly ViTs, but it also presents us with three challenges to address. First, how to effectively reparameterize ViTs with shifts and adds? Previous ShiftAddNet you2020shiftaddnet reparameterizes CNNs by cascading shift layers and add layers, leading to a doubled number of layers or parameters. The customized CUDA kernels of shift and add layers also suffer from much slower training and inference than PyTorch paszke2019pytorch on GPUs. Both motivate a new easy-to-accelerate reparameterization method for ViTs. Second, how to maintain the accuracy after reparameterization? It is expected to see accuracy drops when multiplications are turned into shift and add primitives chen2020addernet ; chen2021empirical . Most works compensate for the accuracy drops by enlarging model sizes leveraging the much improved energy efficiency advantages of shifts and adds you2020shiftaddnet , e.g., up to 196$\times$ unit energy savings than multiplications as shown in Tab. 1. While in ViTs, where input images are split into non-overlapping tokens, one can uniquely leverage the inherent adaptive sensitivity among input tokens. In principle, the essential tokens containing target objects are expected to be processed using more powerful multiplications. In contrast, tokens with less important background information can be handled by cheaper primitives. Such a principle aligns with the spirit of recent token merging bolya2022token and input adaptive methods rao2021dynamicvit ; yin2022avit ; yu2022mia for ViTs. Third, how can we balance the loading and processing time for the aforementioned important and less-important input tokens? For ViTs, using a mixture of multiplication primitives can lead to varying processing speeds for sensitive and non-sensitive tokens. This needs to be balanced; otherwise, intermediate activations may take longer to synchronize before progressing to the next layer.
@@ 2 Related Works
Vision Transformers (ViTs). ViTs vit ; DeiT split images into non-overlapping patches or tokens and have outperformed CNNs on a range of vision tasks. They use a simple encoder-only architecture, primarily made up of attention modules and MLPs. The initial ViTs necessitated costly pre-training on expansive datasets chen2017revisiting , and subsequent advancements like DeiT DeiT and T2T-ViT yuan2021tokens alleviated this by refining training methodologies or crafting innovative tokenization strategies. Later on, a surge of new ViT architecture designs, such as PVT PVT_ICCV , CrossViT chen2021crossvit , PiT heo2021pit , MViT fan2021multiscale , and Swin-Transformer liu2021swin , has emerged to improve the accuracy-efficiency trade-offs using a pyramid-like architecture. Another emergent trend revolves around dynamic ViTs like DynamicViT rao2021dynamicvit , A-ViT yin2022avit , ToME bolya2022token , and MIA-Former yu2022mia , aiming to adaptively eliminate non-essential tokens. Our MoE framework resonates with this concept but addresses those tokens using low-cost multiplication primitives that run concurrently. For edge deployment, solutions such as LeViT graham2021levit , MobileViT mehta2021mobilevit , and EfficientViT cai2022efficientvit ; shen2021efficient employ efficient attention mechanisms or integrate CNN feature extractors. This approach propels ViTs to achieve speeds comparable to CNNs, e.g., MobileNets sandler2018mobilenetv2 . In contrast, our proposed ShiftAddViT is the first to draw from shift&add hardware shortcuts for reparameterizing pre-trained ViTs, boosting end-to-end inference speeds and energy efficiency without hurting accuracy. Therefore, ShiftAddViT is orthogonal to existing ViT designs and can be applied on top of them.
@@
Tremendous efforts have been dedicated to designing efficient ViTs. For instance, to address the costly self-attention module, which possesses quadratic complexity in relation to the number of input tokens, various linear attentions have been introduced. These can be broadly divided into two categories: local attention liu2021swin ; arar2022learned ; tu2022maxvit and kernel-based linear attention katharopoulos2020transformers ; choromanski2021rethinking ; xiong2021nystromformer ; lu2021soft ; cai2022efficientvit ; liu2022ecoformer ; ali2021xcit . For the former category, Swin liu2021swin computes similarity measures among neighboring tokens rather than all tokens. QnA arar2022learned disseminates the attention queries across all tokens. MaxViT tu2022maxvit employs block attention and integrates dilated global attention to capture both local and global information. Regarding the latter category, most methods approximate the softmax function choromanski2021rethinking ; katharopoulos2020transformers ; bolya2022hydra or the complete self-attention matrix xiong2021nystromformer ; lu2021soft using orthogonal features or kernel embeddings, allowing the computational order to switch from $\mathbf{(QK)V}$ to $\mathbf{Q(KV)}$. Furthermore, a few efforts target reducing the multiplication count in ViTs. For instance, Ecoformer liu2022ecoformer introduces a novel binarization paradigm through kernelized hashing for $\mathbf{Q/K}$, ensuring that MatMuls in attentions become solely accumulations. Adder Attention shu2021adder examines the challenges of integrating adder operations into attentions and suggests including an auxiliary identity mapping. ShiftViT wang2022shift incorporates spatial shift operations into attention modules. Different from previous works, our ShiftAddViT focuses on not only attention but also MLPs towards end-to-end ViT speedups. All MatMuls can be transformed into additions/accumulations using quantization, eliminating the need for complex kernelized hashing (KSH) as in liu2022ecoformer . For the remaining linear layers in the attention modules or MLPs, they are converted to bitwise shifts or MoEs, rather than employing spatial shifts like wang2022shift for attentions.
@@
Multiplication-less NNs. Numerous efforts aim to minimize the dense multiplications, which are primary contributors to the time and energy consumption in CNNs and ViTs. In the realm of CNNs, binary networks courbariaux2016binarized ; juefei2017local binarize both activations and weights; AdderNet chen2020addernet ; adder_distillation ; adder_hardware fully replaces multiplications with additions with a commendable accuracy drop; Shift-based networks use either spatial shifts wu2018shift or bitwise shifts elhoushi2021deepshift to replace multiplications. Recently, techniques initially developed for CNNs have found applications in ViTs. For instance, BiTs liu2022bit ; he2022bivit introduce binary transformers that maintain commendable accuracy; Shu et al. and Wang et al. also migrate the add or shift idea to ViTs with the scope limited to attentions shu2021adder ; wang2022shift . Apart from network designs, there are also neural architecture search efforts made towards both accurate and efficient networks you2022shiftaddnas ; li2022searching . As compared to the most related work, i.e., ShiftAddNet you2020shiftaddnet and ShiftAddNAS you2022shiftaddnas , our ShiftAddViT for the first time enables shift&add idea in ViTs without compromising the model accuracy, featuring three more characteristics that make it more suitable for practical applications and deployment: (1) starting from pre-trained ViTs to avoid the costly and tedious training from scratch; (2) focusing on optimization for GPUs rather than on dedicated hardware such as FPGAs/ASICs; (3) seamlessly compatible with the MoE framework to enable switching between the mixture of multiplication primitives, such token-level routing and parallelism are uniquely applicable and designed for ViTs.
@@ 3 Preliminaries
Self-attention and Vision Transformers. Self-attention is a core ingredient of Transformers transformer ; vit , and usually contains multiple heads H with each measuring pairwise correlations among all input tokens to capture global-context information. It can be defined as below:
@@
Shift and Add Primitives. Direct hardware implementation of multiplications is inefficient. Shift and add primitives serve as “shortcuts” for streamlined hardware designs. Shift operations, equivalent to multiplying by powers of two, offer significant savings. As shown in Tab. 1, such shifts can reduce energy use by up to 23.8$\times$ and chip area by 22.3$\times$ compared to multiplications lian2019high ; you2020shiftaddnet , given the INT32 data format and 45nm CMOS technology. Similarly, add operations, another efficient primitive, can achieve up to 31.0$\times$ energy and 25.5$\times$ area savings relative to multiplications. Both primitives have driven numerous network architecture innovations chen2020addernet ; elhoushi2021deepshift ; you2020shiftaddnet ; you2022shiftaddnas .
@@ 4 The Proposed ShiftAddViT
Overview. Given the widespread availability of pre-trained ViT model weights, we aim to fine-tune and transform them into multiplication-reduced ShiftAddViTs in order to decrease runtime latency and enhance energy efficiency. To achieve this, we must reparameterize both the attentions and MLPs in ViTs to shift and add operations. Instead of using cascaded shift layers and add layers as in ShiftAddNet you2020shiftaddnet or designing specific forward and backpropagation schemes like AdderNet chen2020addernet ; shu2021adder , we lean on the simple yet effective approach of ViTs’ long-dependency modeling capability. We propose replacing the dominant MLPs/Linears/MatMuls with shifts and adds, leaving the attention intact. This approach permits ShiftAddViTs to be adapted from existing ViTs without the need for training from scratch. As illustrated in Fig. 1, for attentions, we transform four Linear layers and two MatMuls into shift and add layers, respectively. For MLPs, direct replacement with shift layers results in significant accuracy degradation. Hence, we consider designing a dedicated MoE framework incorporating a blend of multiplication primitives, such as multiplication and shift, to achieve both accurate and efficient final models. These steps transition pre-trained ViTs into ShiftAddViTs, achieving much reduced runtime latency while maintaining a comparable or even higher accuracy.
@@ 4.1 ShiftAddViT: Reparameterization Towards Multiplication-reduced Networks
Reparameterization of Attention. There are two categories of layers in attention modules: MatMuls and Linears, that can be converted to Add and Shift layers, respectively. In order to reparameterize MatMuls to Add layers, we consider performing binary quantization on one operand during MatMuls, e.g., $\mathbf{Q}$ or $\mathbf{K}$ for the MatMul of $\mathbf{QK}$, so that the multiply-and-accumulate (MAC) operations between two matrices will be replaced with merely energy-efficient add operations liu2022ecoformer . Furthermore, to build ShiftAddViT on top of efficient linear attentions wang2020linformer ; liu2022ecoformer ; lu2021soft , we exchange the order of MatMuls from $\mathbf{(QK)V}$ to $\mathbf{Q(KV)}$ for achieving linear complexity w.r.t. the number of input tokens. In this way, the binary quantization will be applied to $\mathbf{K}$ and $\mathbf{Q}$ as illustrated in Fig. 1 (b), leaving the more sensitive $\mathbf{V}$ branch as high precision. This also allows us to insert a lightweight depthwise convolution (DWConv) to $\mathbf{V}$ branch in parallel to enhance local feature capturing capability with negligible overhead ($<$1% of total MACs) following recent SOTA linear attention designs xiong2021nystromformer ; you2023castling . On the other hand, for reparameterizing the left four Linears in the attention module with Shift layers, we resort to sign flips and power-of-two shift parameters to represent the Linear weights elhoushi2021deepshift ; you2020shiftaddnet .
@@
where $\mathbf{X}$ refers to input activations, $\mathbf{W_{S}}$ and $\mathbf{W_{A}}$ are weights in Shift and Add layers, respectively. For Shift weights $\mathbf{W_{S}}=\mathbf{s}\cdot 2^{\mathbf{P}}$, $\mathbf{s}=sign(\mathbf{W})\in\{-1,1\}$ denotes sign flips, $\mathbf{P}=round(log_{2}(abs(\mathbf{W})))$, and $2^{\mathbf{P}}$ represents bitwise shift towards left ($\mathbf{P}>0$) or right ($\mathbf{P}<0$). Both $\mathbf{s}$ and $\mathbf{P}$ are trainable during the training. For Add weights, we have applied binary quantization and will skip zeros directly, the remaining weights correspond to additions or accumulations. Combining those two weak players, i.e., Shift and Add, leads to higher representation capability as also stated in ShiftAddNet you2020shiftaddnet , we highlight that our ShiftAddViT seamlessly achieves such combination for ViTs, without the necessity of cascading both of them as a substitute for one convolutional layer.
Reparameterization of MLPs. MLPs are often overlooked but also dominate ViTs’ runtime latency apart from attentions, especially for the small or medium amount of tokens fan2022adaptable ; zhou2022energon ; qu2022dota . One natural thought is to reparameterize all MLPs with Shift layers as defined above. However, such aggressive design indispensably leads to severe accuracy drops, e.g., $\downarrow$1.18for PVTv1-T PVT_ICCV , (DWConv used between two MLPs in PVTv2 are kept). We resolve this by proposing a new MoE framework to achieve speedups with comparable accuracy, as shown in Fig. 1 (c) and will be elaborated in Sec. 4.2.
@@
  Components Apply Acc. (%) of PVT PVT_ICCV ; wang2022pvt

### v6

@@ 1 Introduction
Vision Transformers (ViTs) have emerged as powerful vision backbone substitutes to convolutional neural networks (CNNs) due to their impressive performance [57, 16]. However, ViTs’ state-of-the-art (SOTA) accuracy comes at the price of prohibitive hardware latency and energy consumption for both training and inference, limiting their prevailing deployment and applications [34, 51], even though pre-trained ViT models are publicly available. The bottlenecks are twofold based on many profiling results: attention and MLPs, e.g., they account for 36%/63% of the FLOPs and 45%/46% of the latency when running DeiT-Base [55] on an NVIDIA Jetson TX2 GPU [72, 19]). Previous efficient ViT solutions mostly focus on macro-architecture designs [21, 39, 32, 55, 9, 61] and linear attention optimization [59, 50, 5, 13, 6, 73]. However, they pay less attention on reducing the underlying dominant multiplications as well as on co-optimizing both attention and MLPs.
We identify that one of the most effective yet still missing opportunities for accelerating ViTs is to reparameterize their redundant multiplications with a mixture of multiplication primitives, i.e., bitwise shifts and adds. The idea is drawn from the common hardware design practice in computer architecture or digital signal processing. That is, the multiplication can be replaced by bitwise shifts and adds [67, 22]. Such a hardware-inspired approach can lead to an efficient and fast hardware implementation without compromising the model accuracy. As such, this paper embodies the shift&add idea in ViTs towards a new type of multiplication-reduced model. Moreover, unlike previous ShiftAddNet [70] which requires full and slow training and dedicated hardware accelerator support, this work starts from pre-trained ViTs to avoid the tremendous cost of training from scratch and targets end-to-end inference speedups, i.e., accelerating both attention and MLPs, on GPUs.
@@
The aforementioned shift&add concept uniquely inspires our design of naturally hardware-friendly ViTs, but it also presents us with three challenges to address. First, how to effectively reparameterize ViTs with shifts and adds? Previous ShiftAddNet [70] reparameterizes CNNs by cascading shift layers and add layers, leading to a doubled number of layers or parameters. The customized CUDA kernels of shift and add layers also suffer from much slower training and inference than PyTorch [42] on GPUs. Both motivate a new easy-to-accelerate reparameterization method for ViTs. Second, how to maintain the accuracy after reparameterization? It is expected to see accuracy drops when multiplications are turned into shift and add primitives [8, 11]. Most works compensate for the accuracy drops by enlarging model sizes leveraging the much improved energy efficiency advantages of shifts and adds [70], e.g., up to 196$\times$ unit energy savings than multiplications as shown in Tab. 1. While in ViTs, where input images are split into non-overlapping tokens, one can uniquely leverage the inherent adaptive sensitivity among input tokens. In principle, the essential tokens containing target objects are expected to be processed using more powerful multiplications. In contrast, tokens with less important background information can be handled by cheaper primitives. Such a principle aligns with the spirit of recent token merging [3] and input adaptive methods [45, 68, 74] for ViTs. Third, how can we balance the loading and processing time for the aforementioned important and less-important input tokens? For ViTs, using a mixture of multiplication primitives can lead to varying processing speeds for sensitive and non-sensitive tokens. This needs to be balanced; otherwise, intermediate activations may take longer to synchronize before progressing to the next layer.
@@ 2 Related Works
Vision Transformers (ViTs). ViTs [16, 55] split images into non-overlapping patches or tokens and have outperformed CNNs on a range of vision tasks. They use a simple encoder-only architecture, primarily made up of attention modules and MLPs. The initial ViTs necessitated costly pre-training on expansive datasets [52], and subsequent advancements like DeiT [55] and T2T-ViT [75] alleviated this by refining training methodologies or crafting innovative tokenization strategies. Later on, a surge of new ViT architecture designs, such as PVT [60], CrossViT [7], PiT [26], MViT [18], and Swin-Transformer [35], has emerged to improve the accuracy-efficiency trade-offs using a pyramid-like architecture. Another emergent trend revolves around dynamic ViTs like DynamicViT [45], A-ViT [68], ToME [3], and MIA-Former [74], aiming to adaptively eliminate non-essential tokens. Our MoE framework resonates with this concept but addresses those tokens using low-cost multiplication primitives that run concurrently. For edge deployment, solutions such as LeViT [21], MobileViT [39], and EfficientViT [5, 50] employ efficient attention mechanisms or integrate CNN feature extractors. This approach propels ViTs to achieve speeds comparable to CNNs, e.g., MobileNets [47]. In contrast, our proposed ShiftAddViT is the first to draw from shift&add hardware shortcuts for reparameterizing pre-trained ViTs, boosting end-to-end inference speeds and energy efficiency without hurting accuracy. Therefore, ShiftAddViT is orthogonal to existing ViT designs and can be applied on top of them.
@@
Tremendous efforts have been dedicated to designing efficient ViTs. For instance, to address the costly self-attention module, which possesses quadratic complexity in relation to the number of input tokens, various linear attentions have been introduced. These can be broadly divided into two categories: local attention [35, 2, 56] and kernel-based linear attention [30, 13, 65, 38, 5, 34, 1]. For the former category, Swin [35] computes similarity measures among neighboring tokens rather than all tokens. QnA [2] disseminates the attention queries across all tokens. MaxViT [56] employs block attention and integrates dilated global attention to capture both local and global information. Regarding the latter category, most methods approximate the softmax function [13, 30, 4] or the complete self-attention matrix [65, 38] using orthogonal features or kernel embeddings, allowing the computational order to switch from $\mathbf{(QK)V}$ to $\mathbf{Q(KV)}$. Furthermore, a few efforts target reducing the multiplication count in ViTs. For instance, Ecoformer [34] introduces a novel binarization paradigm through kernelized hashing for $\mathbf{Q/K}$, ensuring that MatMuls in attentions become solely accumulations. Adder Attention [51] examines the challenges of integrating adder operations into attentions and suggests including an auxiliary identity mapping. ShiftViT [58] incorporates spatial shift operations into attention modules. Different from previous works, our ShiftAddViT focuses on not only attention but also MLPs towards end-to-end ViT speedups. All MatMuls can be transformed into additions/accumulations using quantization, eliminating the need for complex kernelized hashing (KSH) as in [34]. For the remaining linear layers in the attention modules or MLPs, they are converted to bitwise shifts or MoEs, rather than employing spatial shifts like [58] for attentions.
@@
Multiplication-less NNs. Numerous efforts aim to minimize the dense multiplications, which are primary contributors to the time and energy consumption in CNNs and ViTs. In the realm of CNNs, binary networks [14, 29] binarize both activations and weights; AdderNet [8, 66, 62] fully replaces multiplications with additions with a commendable accuracy drop; Shift-based networks use either spatial shifts [64] or bitwise shifts [17] to replace multiplications. Recently, techniques initially developed for CNNs have found applications in ViTs. For instance, BiTs [36, 25] introduce binary transformers that maintain commendable accuracy; Shu et al. and Wang et al. also migrate the add or shift idea to ViTs with the scope limited to attentions [51, 58]. Apart from network designs, there are also neural architecture search efforts made towards both accurate and efficient networks [71, 31]. As compared to the most related work, i.e., ShiftAddNet [70] and ShiftAddNAS [71], our ShiftAddViT for the first time enables shift&add idea in ViTs without compromising the model accuracy, featuring three more characteristics that make it more suitable for practical applications and deployment: (1) starting from pre-trained ViTs to avoid the costly and tedious training from scratch; (2) focusing on optimization for GPUs rather than on dedicated hardware such as FPGAs/ASICs; (3) seamlessly compatible with the MoE framework to enable switching between the mixture of multiplication primitives, such token-level routing and parallelism are uniquely applicable and designed for ViTs.
@@ 3 Preliminaries
Self-attention and Vision Transformers. Self-attention is a core ingredient of Transformers [57, 16], and usually contains multiple heads H with each measuring pairwise correlations among all input tokens to capture global-context information. It can be defined as below:
@@
Shift and Add Primitives. Direct hardware implementation of multiplications is inefficient. Shift and add primitives serve as “shortcuts” for streamlined hardware designs. Shift operations, equivalent to multiplying by powers of two, offer significant savings. As shown in Tab. 1, such shifts can reduce energy use by up to 23.8$\times$ and chip area by 22.3$\times$ compared to multiplications [33, 70], given the INT32 data format and 45nm CMOS technology. Similarly, add operations, another efficient primitive, can achieve up to 31.0$\times$ energy and 25.5$\times$ area savings relative to multiplications. Both primitives have driven numerous network architecture innovations [8, 17, 70, 71].
@@ 4 The Proposed ShiftAddViT
Overview. Given the widespread availability of pre-trained ViT model weights, we aim to fine-tune and transform them into multiplication-reduced ShiftAddViTs in order to decrease runtime latency and enhance energy efficiency. To achieve this, we must reparameterize both the attentions and MLPs in ViTs to shift and add operations. Instead of using cascaded shift layers and add layers as in ShiftAddNet [70] or designing specific forward and backpropagation schemes like AdderNet [8, 51], we lean on the simple yet effective approach of ViTs’ long-dependency modeling capability. We propose replacing the dominant MLPs/Linears/MatMuls with shifts and adds, leaving the attention intact. This approach permits ShiftAddViTs to be adapted from existing ViTs without the need for training from scratch. As illustrated in Fig. 1, for attentions, we transform four Linear layers and two MatMuls into shift and add layers, respectively. For MLPs, direct replacement with shift layers results in significant accuracy degradation. Hence, we consider designing a dedicated MoE framework incorporating a blend of multiplication primitives, such as multiplication and shift, to achieve both accurate and efficient final models. These steps transition pre-trained ViTs into ShiftAddViTs, achieving much reduced runtime latency while maintaining a comparable or even higher accuracy.
@@ 4.1 ShiftAddViT: Reparameterization Towards Multiplication-reduced Networks
Reparameterization of Attention. There are two categories of layers in attention modules: MatMuls and Linears, that can be converted to Add and Shift layers, respectively. In order to reparameterize MatMuls to Add layers, we consider performing binary quantization on one operand during MatMuls, e.g., $\mathbf{Q}$ or $\mathbf{K}$ for the MatMul of $\mathbf{QK}$, so that the multiply-and-accumulate (MAC) operations between two matrices will be replaced with merely energy-efficient add operations [34]. Furthermore, to build ShiftAddViT on top of efficient linear attentions [59, 34, 38], we exchange the order of MatMuls from $\mathbf{(QK)V}$ to $\mathbf{Q(KV)}$ for achieving linear complexity w.r.t. the number of input tokens. In this way, the binary quantization will be applied to $\mathbf{K}$ and $\mathbf{Q}$ as illustrated in Fig. 1 (b), leaving the more sensitive $\mathbf{V}$ branch as high precision. This also allows us to insert a lightweight depthwise convolution (DWConv) to $\mathbf{V}$ branch in parallel to enhance local feature capturing capability with negligible overhead ($<$1% of total MACs) following recent SOTA linear attention designs [65, 73]. On the other hand, for reparameterizing the left four Linears in the attention module with Shift layers, we resort to sign flips and power-of-two shift parameters to represent the Linear weights [17, 70].
@@
where $\mathbf{X}$ refers to input activations, $\mathbf{W_{S}}$ and $\mathbf{W_{A}}$ are weights in Shift and Add layers, respectively. For Shift weights $\mathbf{W_{S}}=\mathbf{s}\cdot 2^{\mathbf{P}}$, $\mathbf{s}=sign(\mathbf{W})\in\{-1,1\}$ denotes sign flips, $\mathbf{P}=round(log_{2}(abs(\mathbf{W})))$, and $2^{\mathbf{P}}$ represents bitwise shift towards left ($\mathbf{P}>0$) or right ($\mathbf{P}<0$). Both $\mathbf{s}$ and $\mathbf{P}$ are trainable during the training. For Add weights, we have applied binary quantization and will skip zeros directly, the remaining weights correspond to additions or accumulations. Combining those two weak players, i.e., Shift and Add, leads to higher representation capability as also stated in ShiftAddNet [70], we highlight that our ShiftAddViT seamlessly achieves such combination for ViTs, without the necessity of cascading both of them as a substitute for one convolutional layer.
Reparameterization of MLPs. MLPs are often overlooked but also dominate ViTs’ runtime latency apart from attentions, especially for the small or medium amount of tokens [19, 78, 43]. One natural thought is to reparameterize all MLPs with Shift layers as defined above. However, such aggressive design indispensably leads to severe accuracy drops, e.g., $\downarrow$1.18for PVTv1-T [60], (DWConv used between two MLPs in PVTv2 are kept). We resolve this by proposing a new MoE framework to achieve speedups with comparable accuracy, as shown in Fig. 1 (c) and will be elaborated in Sec. 4.2.
@@
  Components Apply Acc. (%) of PVT [60, 61]